OpenStudy (anonymous):

If x - 3y = 0 and 5x - y = 14, then 6x - 4y =

OpenStudy (psymon):

If x - 3y = 0, then this means x = 3y. Replace x in your second equation with 3y and you'll then be able to solve for y, which will then help you get the full answer.

OpenStudy (anonymous):

Stack those 2 equations on top of one another and add 'em together x - 3y = 0 5x - y = 14 ----------------- 6x - 4y = 14 Answer 14

OpenStudy (psymon):

Ehh, all he does is give answers -_-